# A Photon Cloud Induced from an Axion Cloud

### v1

@@ III A Photon Cloud Induced from an axion cloud
Next, we try to solve the first-order induced EM field $F^{\mu\nu}_{(1)}$ in the Newman-Penrose (NP) formalism, in which using the Kinnersley’s null tetrad [30]
@@
where $\varphi_{0}$ is the ingoing radiation with spin weight $s=1$ and $\varphi_{2}$ represents outgoing radiation with spin weight $s=-1$. Note that there are only two real degree of freedoms though the three components are complex. The EM tensor can be reconstructed as
@@
In fact, in the scenario we are considering, the axion cloud is the source for the induced EM field that
@@
$\phi_{(0)}=\alpha_{1}^{2}\phi_{0}\frac{r}{M}e^{-\frac{r}{2M}\alpha_{1}^{2}}\cos{\left(\varphi-\omega_{R}t\right)}\sin{\theta}~{},$ (27)
which is valid for $\alpha\lesssim 0.2$ even at large black hole spin, and $\omega_{R}\simeq\mu_{1}$ is the oscillating frequency.
It can be separated into two parts
@@
with $R(r)=\alpha_{1}^{2}\phi_{0}\frac{r}{2M}e^{-\frac{r}{2M}\alpha_{1}^{2}}$ and $S(\theta)=\sin{\theta}$.
@@
we can obtain the three decoupled differential equations respectively for $f_{01}\left(r,\theta\right)$, $f_{11}\left(r,\theta\right)$ and $f_{21}\left(r,\theta\right)$ after dropping off the time $t$ and azimuthal angle $\varphi$ dependence by $\omega_{1}=\omega$ and $m_{1}=m=1$. We can also do that for $\phi_{2}$.
@@
We replace $\omega=\mu_{1}=\alpha_{1}/M$, $a_{0}\equiv a/M$, $r_{0}\equiv r/M$, and define new dimensionless quantities $f_{01M}\equiv Mf_{01}\left(r,\theta\right)$, $f_{11M}\equiv Mf_{11}\left(r,\theta\right)$, $f_{\rm 21M}\equiv Mf_{21}\left(r,\theta\right)$. After absorbing some coefficients from the source part into the coupling constant $k_{a}$, we now use a new dimensionless coupling constant $\varepsilon\equiv\alpha_{1}^{2}B_{0}M\phi_{0}k_{a}$ as the small quantity for expansion, then the new field equations are dimensionless
${\rm eq}f_{\rm 01M}:\left\{\frac{\partial^{2}f_{\rm 01M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 01M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 01M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 01M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 01M}\left(r_{0},\theta\right),a_{0},\alpha_{1}\right\}~{},$ (30)
${\rm eq}f_{\rm 11M}:\left\{\frac{\partial^{2}f_{\rm 11M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 11M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 11M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 11M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 11M}\left(r_{0},\theta\right),a_{0},\alpha_{1}\right\}~{},$ (31)
${\rm eq}f_{\rm 21M}:\left\{\frac{\partial^{2}f_{\rm 21M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 21M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 21M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 21M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 21M}\left(r_{0},\theta\right),a_{0},\alpha_{1}\right\}~{},$ (32)
and there are only two dimensionless parameters $a_{0}$ and $\alpha_{1}$.
@@
Different from the quasi-normal modes and bound states, here the frequency of the induced EM field is already fixed by the frequency of the source (axion cloud), it is not an eigenvalue problem and no boundary condition should be imposed. Actually we can obtain the asymptotic behaviors of the functions at the boundaries by expanding the equations (we choose decaying solutions at spatial infinity), and then solve the equations numerically.
$f_{\rm 01M}\left(r_{\rm h0},\theta\right)=\frac{2\sqrt{2}a_{0}r_{h0}^{3}e^{-\alpha_{1}^{2}r_{\rm h0}/2}\sin{\theta}\sin{2\theta}}{\left(2\alpha_{1}r_{h0}-a_{0}-2i\sqrt{1-a_{0}^{2}}\right)\left(r_{\rm h0}^{2}+a_{0}^{2}\cos^{2}{\theta}\right)^{2}}~{},$ (33)
$f_{\rm 11M}\left(r_{\rm h0},\theta\right)=\frac{2i\sqrt{1-a_{0}^{2}}r_{h0}^{3}e^{-\alpha_{1}^{2}r_{\rm h0}/2}\sin{2\theta}}{\left(a_{0}-2\alpha_{1}r_{h0}\right)\left(r_{\rm h0}^{2}+a_{0}^{2}\cos^{2}{\theta}\right)^{2}}~{},$ (34)
$f_{\rm 21M}\left(r_{\rm h0},\theta\right)=0~{},$ (35)
$f_{\rm 01M}\left(r_{0}\to\infty,\theta\right)=-\frac{3i\sin^{2}{\theta}e^{-\alpha_{1}^{2}r_{0}/2}}{\sqrt{2}\alpha_{1}}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha_{1}^{2}r_{0}/2}\right)~{},$ (36)
$f_{\rm 11M}\left(r_{0}\to\infty,\theta\right)=-\frac{(\alpha_{1}+3i)\sin{2\theta}e^{-\alpha_{1}^{2}r_{0}/2}}{4\alpha_{1}}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha_{1}^{2}r_{0}/2}\right)~{},$ (37)
$f_{\rm 21M}\left(r_{0}\to\infty,\theta\right)=-\frac{3i\sin^{2}{\theta}e^{-\alpha_{1}^{2}r_{0}/2}}{2\sqrt{2}\alpha_{1}}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha_{1}^{2}r_{0}/2}\right)~{},$ (38)
In addition, it is clear that the asymptotic behaviors for $f_{\rm 01M}$ and $f_{\rm 11M}$ at event horizon $r_{\rm h0}$ are only valid when $\alpha_{1}\neq\frac{a_{0}}{2r_{\rm h0}}$, which is exactly the critical condition for whether superradiance occurs. In this critical case with $\alpha_{1}=\frac{a_{0}}{2r_{\rm h0}}$, the imaginary part $\omega_{I}$ disappears in the original field equation for the axion field, so that the whole frequency is not complex and the original boundary condition at event horizon is not valid, in which case another regular condition at origin should be applied. Here we do not go further on this case.
Similarly, for the NP scalars induced from $\phi_{2}$
@@
which means that with the above transformations, we can start from one set of the equations to the other set. These symmetries from analysis are consistent with the numerical results, thus here we only exhibit the distribution for those from $\phi_{1}$ in FIG. 1.
Figure 1: The real and imaginal parts of $f_{01M}\left(r_{0},\theta\right)$, $f_{11M}\left(r_{0},\theta\right)$, $f_{21M}\left(r_{0},\theta\right)$ induced from $\phi_{1}$ are plotted with $a_{0}=0.5$ and $\alpha_{1}=0.1$ in the scale of $200M$, where the superradiant condition is satisfied $\alpha_{1}<\frac{a_{0}}{2r_{\rm h0}}$.
From FIG. 1 we can see that, the ingoing radiation $f_{01M}$ is slightly larger than the outgoing radiation $f_{21M}$, and the residential part $f_{11M}$ presents a ripple-like pattern. Usually the residential $\varphi_{1}$ part is neglected, since its equation is not separable even without any source and it is not related to ingoing or outgoing part. But here, it is shown that the imaginary part of $f_{11M}$ gets much larger amplitude than the other parts, indicating that the residential part is the primary constituent of the EM photon cloud we are studying. In fact, both $\varphi_{0}$ and $\varphi_{2}$ contain ingoing and outgoing parts, people assign ingoing/outgoing radiation to $\varphi_{0}$/$\varphi_{2}$ due to which one is dominant at spatial infinity.
Subsequently, we can construct the EM tensor induced from $\phi_{1}$ using (19)
$F_{\mu\nu}^{1}=4\cos{\left(m\varphi-\omega t\right)}{\rm Re}\left[P_{\mu\nu}^{1}\left(r,\theta\right)\right]-4\sin{\left(m\varphi-\omega t\right)}{\rm Im}\left[P_{\mu\nu}^{1}\left(r,\theta\right)\right]~{},$ (45)
$P_{\mu\nu}^{1}\left(r,\theta\right)\equiv f_{01}\left(r,\theta\right)\overline{m}_{[\mu}n_{\nu]}+f_{11}\left(r,\theta\right)n_{[\mu}l_{\nu]}+f_{11}\left(r,\theta\right)m_{[\mu}\overline{m}_{\nu]}+f_{21}\left(r,\theta\right)l_{[\mu}m_{\nu]}~{}.$ (46)
Similarly, for the EM tensor induced from $\phi_{2}$ we have
$F_{\mu\nu}^{2}=4\cos{\left(m\varphi-\omega t\right)}{\rm Re}\left[P_{\mu\nu}^{2}\left(r,\theta\right)\right]+4\sin{\left(m\varphi-\omega t\right)}{\rm Im}\left[P_{\mu\nu}^{2}\left(r,\theta\right)\right]~{},$ (47)
$P_{\mu\nu}^{2}\left(r,\theta\right)\equiv f_{02}\left(r,\theta\right)\overline{m}_{[\mu}n_{\nu]}+f_{12}\left(r,\theta\right)n_{[\mu}l_{\nu]}+f_{12}\left(r,\theta\right)m_{[\mu}\overline{m}_{\nu]}+f_{22}\left(r,\theta\right)l_{[\mu}m_{\nu]}~{},$ (48)
then the whole EM field induced from the axion cloud $\phi_{(0)}=\phi_{1}+\phi_{2}$ evolves as
$F_{\mu\nu}^{(1)}=F_{\mu\nu}^{1}+F_{\mu\nu}^{2}=\cos{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}\left(r,\theta\right)+\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}\left(r,\theta\right)~{},$ (49)
@@
So far we have considered an EM photon cloud sourced by an axion cloud surrounding a Kerr black hole which is described by (49). This is one of the main results of our work. Moreover, it means that for an arbitrary current $J^{a}=e^{\pm i\omega t\mp im\varphi}S\left(r,\theta\right)$ as the source we will always get a photon cloud evolving as
$F_{\mu\nu}=4\cos{\left(m\varphi-\omega t\right)}{\rm Re}\left[P_{\mu\nu}\left(r,\theta\right)\right]\pm 4\sin{\left(m\varphi-\omega t\right)}{\rm Im}\left[P_{\mu\nu}\left(r,\theta\right)\right]$.
In fact, even for photons without any couplings, they can have an effective mass given by plasma and make superradiance to happen [32, 33].

### v4

@@ III A Photon Cloud Induced from an axion cloud
In this Section we will solve the first-order induced EM field $F^{\mu\nu}_{(1)}$ in the Newman-Penrose (NP) formalism, in which using the Kinnersley’s null tetrad [44]
@@
where $\varphi_{0}$ is the ingoing radiation with spin weight $s=1$ and $\varphi_{2}$ represents outgoing radiation with spin weight s=1s=-1. Note that there are only two real degree of freedoms though the three components are complex. The EM tensor can be reconstructed as
@@
In fact, in our framework, the axion cloud is the source for the induced EM field
@@
$\phi_{(0)}=\alpha^{2}\phi_{0}\frac{r}{M}e^{-\frac{r}{2M}\alpha^{2}}\cos{\left(\varphi-\omega_{R}t\right)}\sin{\theta}~,$ (33)
which is valid for $\alpha\lesssim 0.2$ even at large black hole spin, and $\omega_{R}\simeq\mu$ is the oscillating frequency. It can be separated into two parts
@@
with $R(r)=\alpha^{2}\phi_{0}\frac{r}{2M}e^{-\frac{r}{2M}\alpha^{2}}$ and $S(\theta)=\sin{\theta}$.
@@
we can obtain the three decoupled differential equations respectively for $f_{01}\left(r,\theta\right)$, $f_{11}\left(r,\theta\right)$ and $f_{21}\left(r,\theta\right)$ after dropping off the time $t$ and azimuthal angle $\varphi$ dependence by $\omega_{1}=\omega$ and $m_{1}=m=1$. We can also do that for $\phi_{2}$. Then we replace $\omega=\mu=\alpha/M$, $a_{0}\equiv a/M$, $r_{0}\equiv r/M$, and define new dimensionless quantities $f_{01M}\equiv Mf_{01}\left(r,\theta\right)$, $f_{11M}\equiv Mf_{11}\left(r,\theta\right)$, $f_{\rm 21M}\equiv Mf_{21}\left(r,\theta\right)$. After absorbing some coefficients from the source part into the coupling constant $k_{a}$, we now use a new dimensionless coupling constant $\varepsilon\equiv\alpha^{2}B_{0}M\phi_{0}k_{a}$ as the small quantity for expansion, then the new field equations are dimensionless
${\rm eq}f_{\rm 01M}:\left\{\frac{\partial^{2}f_{\rm 01M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 01M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 01M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 01M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 01M}\left(r_{0},\theta\right),a_{0},\alpha\right\}~,$ (36)
${\rm eq}f_{\rm 11M}:\left\{\frac{\partial^{2}f_{\rm 11M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 11M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 11M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 11M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 11M}\left(r_{0},\theta\right),a_{0},\alpha\right\}~,$ (37)
${\rm eq}f_{\rm 21M}:\left\{\frac{\partial^{2}f_{\rm 21M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 21M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 21M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 21M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 21M}\left(r_{0},\theta\right),a_{0},\alpha\right\}~,$ (38)
and there are only two dimensionless parameters $a_{0}$ and $\alpha$.
@@
Different from the quasi-normal modes and bound states, here the frequency of the induced EM field is already fixed by the frequency of the source (axion cloud), it is not an eigenvalue problem and no boundary condition should be imposed. Actually we can obtain the asymptotic behaviors of the functions at the boundaries by expanding the equations (we choose decaying solutions at spatial infinity), and then solve the equations numerically. The asymptotic behaviors of $f_{01M}$, $f_{11M}$ and $f_{\rm 21M}$ at the boundaries can be found in Appendix A. Similarly, for the NP scalars induced from $\phi_{2}$
@@
which means that with the above transformations, we can start from one set of the equations to the other set. These symmetries from analysis are consistent with the numerical results, thus here we only exhibit the distribution for those from $\phi_{1}$ in FIG. 1, where the axes $x=r_{0}\sin{\theta}$ and $y=r_{0}\cos{\theta}$ are defined to describe the cross section.
Figure 1: The real and imaginal parts of $f_{01M}\left(r_{0},\theta\right)$, $f_{11M}\left(r_{0},\theta\right)$, $f_{21M}\left(r_{0},\theta\right)$ induced from $\phi_{1}$ are plotted with $a_{0}=0.5$ and $\alpha=0.1$ in the scale of $200M$, where the superradiant condition is satisfied $\alpha<\frac{a_{0}}{2r_{\rm h0}}$.
From FIG. 1 we can see that, the ingoing radiation $f_{01M}$ is slightly larger than the outgoing radiation $f_{21M}$, and the residential part $f_{11M}$ presents a ripple-like pattern. Usually the residential $\varphi_{1}$ part is neglected, since its equation is not separable even without any source and it is not related to ingoing or outgoing part. But here, it is shown that the imaginary part of $f_{11M}$ gets much larger amplitude than the other parts, indicating that the residential part is the primary constituent of the EM photon cloud we are concerned with. In fact, both $\varphi_{0}$ and $\varphi_{2}$ contain ingoing and outgoing parts, people assign ingoing/outgoing radiation to $\varphi_{0}$$\varphi_{2}$ due to which one is dominant at spatial infinity.
Subsequently, we can construct the whole EM tensor induced from $\phi_{1}$ and $\phi_{2}$ as
$F_{\mu\nu}^{(1)}=\cos{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}\left(r,\theta\right)+\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}\left(r,\theta\right)~,$ (43)
where
@@
So far we have considered an EM photon cloud sourced by an axion cloud surrounding a Kerr black hole which is described by (43). This is one of the main results of our work. Moreover, it means that for an arbitrary current $J^{a}=\Sigma e^{\pm i\omega t\mp im\varphi}S\left(r,\theta\right)$ as the source we will always get a photon cloud evolving as (43). In fact, even photons without any couplings can acquire an effective mass in a plasma, allowing superradiance to occur for some primordial black holes [46, 41].
In [31], a detailed study of the coupled axion-Maxwell system was performed on a black hole background, showing that, in the presence of plasmas, the EM instability remains controlled by the axion-photon coupling, even for relatively dense plasmas. In our case, the frequency of the induced photon cloud is twice the EM frequency in their work, making it less susceptible to plasma effects. Moreover, as shown in FIG. 3, the induced EM scalar $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ is concentrated in regions away from both the equatorial plane and the polar axis, which renders it less affected by the accretion disk and relativistic jets.
Importantly, the photon component of the cloud does not correspond to an on-shell propagating photon mode, but instead constitutes part of a mixed axion–photon propagation eigenstate undergoing superradiant amplification. Consequently, the existence of a photon component is not excluded by plasma screening.
Furthermore, we emphasize that although the induced photon cloud is initially bound, it corresponds to unstable bound configurations that can subsequently escape to spatial infinity. Since photons experience no potential barrier at spatial infinity where the contribution from the axion cloud vanishes. From a particle interpretation, the photon component can be regarded as effectively free once it is released from the axion cloud into propagating EM modes.
We now turn to the evolution of the induced EM photon cloud. The parity violation term $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and the contraction of the EM field strength tensor $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ can be obtained as ($m=1$)
which shows that except for the evolving parts with a double frequency $2\omega$, there are also residential parts that are time-independent. The $r,\theta$-dependent components $F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}$, $F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}$, $F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$, $F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}$, $F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}$ and $F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}$ are plotted in FIG. 2-3. Their expressions can be found in Appendix B.
Figure 2: The components $M^{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos},~M^{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin},~M^{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$ of the parity violation term $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ are plotted with $a_{0}=0.8$ and $\alpha=0.1,0.2$.
Figure 3: The components $M^{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos},~M^{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin},~M^{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}$ of the contraction of the EM field strength tensor $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ are plotted with $a_{0}=0.8$ and $\alpha=0.1,0.2$.
We study the effects of the two parameters $a_{0},~\alpha$ and find that the spin parameter $a_{0}$ only has slight impacts on both the amplitude and the structure of the distribution for each components of $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$, while these quantities are sensitive to the fine structure constant $\alpha$. Therefore, in FIG. 2-3 we plot the figures for the components of $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ with fixed $a_{0}=0.8$ and different values of $\alpha$. From the figures we can see that, for all the components, the increase of $\alpha$ leads to a significant decrease in amplitude and a slight concentration on the distribution, which indicates that for fixed black hole mass $M$ the smaller axion mass $\mu$ (larger reduced Compton wavelength) makes the interaction with photons more efficient and stronger. This is consistent with the idea that very light (or massless) particles may have a near-degeneracy with the photons, leading to a mixing phenomenon in which a coherent superposition of the two arises [34]. Also, the slight concentration on the distribution should result from the concentration of the axion cloud with larger $\alpha$, where larger axion mass $\mu$ (smaller reduced Compton wavelength) makes the axion cloud concentrate. In general, the induced EM scalars $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ concentrate in the sine parts $M^{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}$ and $M^{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}$.
It is essential to verify the validity range of the perturbative framework with respect to the EM field, ensuring that $F_{\mu\nu}^{(0)}\gg\varepsilon F_{\mu\nu}^{(1)}$. Using $F_{\mu\nu}^{(0)}\sim B_{0}$, this requirement leads to
$\alpha^{2}\phi_{0}k_{a}MF_{\mu\nu}^{(1)}\ll 1~.$ (48)
For the case with $a_{0}=0.8$ and $\alpha=0.1$, the maximum value of $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ reaches approximately $8\times 10^{4}$, which gives $\phi_{0}k_{a}\ll 0.35$. While for the case with $a_{0}=0.8$ and $\alpha=0.2$, the constraint is significantly relaxed as $\phi_{0}k_{a}\ll 1.25$. Note that the quantity $\phi_{0}k_{a}$ is dimensionless in both geometrized and natural units. Remarkably, the resulting constraint is independent of the background field strength $B_{0}$.